# RLIE: Rule Generation with Logistic Regression, Iterative Refinement, and Evaluation for Large Language Models

### v1

@@ Abstract
Nowadays, Large Language Models (LLMs) are able to propose rules in natural language, overcoming constrains of a predefined predicate space inherent in traditional rule learning. However, existing methods using LLMs often overlook the combination effects of rules, and the potential of coupling LLMs with probabilistic rule learning to ensure robust inference is not fully explored. To address this gap, we introduce RLIE, a unified framework that integrates LLMs with probabilistic modeling to learn a set of probabilistic rules. The RLIE framework comprises four stages: (1) Rule generation, where a LLM proposes and filters candidate rules; (2) Logistic regression, which learns the probabilistic weights of the rules for global selection and calibration; (3) Iterative refinement, which continuously optimizes the rule set based on prediction errors; and (4) Evaluation, which compares the performance of the weighted rule set as a direct classifier against various methods of injecting the rules into an LLM. Generated rules are the evaluated with different inference strategies on multiple real-world datasets. While applying rules directly with corresponding weights brings us superior performance, prompting LLMs with rules, weights and classification results from the logistic model will surprising degrade the performance. This result aligns with the observation that LLMs excel at semantic generation and interpretation but are less reliable at fine-grained, controlled probabilistic integration. Our work investigates the potentials and limitations of using LLMs for inductive reasoning tasks, proposing a unified framework which integrates LLMs with classic probabilistic rule combination methods, paving the way for more reliable neuro-symbolic reasoning systems.
@@ 1 Introduction
In data-driven applications and scientific discovery, the goal is not merely to predict outcomes, but to construct a set of verifiable, reusable, and composable theories(Zhou et al., 2024; Yang et al., 2024a; Minh et al., 2022). These theories can enable explainable, auditable decisions while driving the discovery of new knowledge and underlying structures(Yang et al., 2023; 2024b). These theories can be expressed in formal, structural statements(Cohen et al., 1995; Cropper & Morel, 2021) or natural language hypotheses(Zhou et al., 2024), and they share a common characteristic: they are declarative, testable, and self-contained discriminative patterns that yield predictions verifiable by external evidence [^1] In this paper, we do not distinguish between the terms “rule” and “hypothesis”, and will use “rule” throughout the text for consistency.. The challenge of how to learn and integrate a set of such collaborative patterns to both uncover regularities and advance knowledge accumulation has long been a central problem in machine learning and artificial intelligence(Bazgir et al., 2025).
Take the binary classification task of spam detection as an example. A set of rules might include (1) $\mathrm{HasToken}(\text{`win prize'},x)\land\mathrm{HasAttachment}(x)$; (2) $\mathrm{DomainInBlacklist}(x)\land\mathrm{HasShortLink}(x)$; and (3) $\mathrm{TitleAllCaps}(x)\land\mathrm{HasToken}(\text{`urgent'},x)$, all implying the head “is spam”. These rules cover different facets of the problem, and must work in concert. A classic approach to use the rules to do reasoning is deterministic aggregation, where a positive classification is made, if any rule is satisfied(Fürnkranz & Kliegr, 2015). A probabilistic approach, in contrast, treats each rule’s classification result as a binary indicator ($r_{j}\in\{0,1\}$) and employs a statistical model, such as logistic regression, to learn a global set of weights for combining these signals into a unified prediction(Ruczinski et al., 2003; Friedman & Popescu, 2008). While the well-defined structure of rules offers significant advantages, traditional rule learning methods are constrained by a predefined predicate space, which limits their ability to handle unstructured data and open-ended semantics, thereby restricting their applicability in broader scenarios(Cerna & Cropper, 2024; Hocquette et al., 2024).
The recent advent of Large Language Models (LLMs) presents a new opportunity for rule learning. Given a few examples and task context, LLMs can directly generate rules in natural language(Ellis, 2023; Singh et al., 2022). Their linguistic richness and contextual alignment facilitate the articulation of fine-grained conditions, relations and exceptions, covering patterns that are difficult to capture with manually defined predicates. However, existing methods often focus on the iterative optimization of a single rule(Qiu et al., 2023) or the independent generation and simple aggregation of multiple rules(Yang et al., 2023; Zhou et al., 2024; Yang et al., 2024b), failing to jointly learn and compress the rules as a cohesive set. Concurrently, probabilistic approaches in the context of LLMs remains largely unexplored, as LLMs’ ability to robustly learn the rule weights from data and apply them to perform reliable probabilistic reasoning during inference is not systematically examined yet. This gap motivates our pursuit of a framework that combines the generative and expressive capabilities of LLMs with the probabilistic composition and calibration of traditional methods to produce a flexible, robust and reusable rule set.
To achieve this synergy, we propose RLIE, a unified framework for LLM-based rule learning that integrates Rule generation, Logistic regression, Iterative refinement and Evaluation. First, in the rule generation stage, an LLM produces a set of independent, individually assessable natural language rules based on a few examples and the task context. Second, for probabilistic combination, a regularized logistic regression model is employed to perform global weighting and selection on the rules, yielding a compact, interpretable rule set with importance weights. Third, in the iterative refinement stage, challenging examples are identified based on the prediction errors. These examples, along with the current rule set, are fed back to the LLM to prompt reflective generation of improved rules until convergence. Finally, a evaluation of two primary inference strategies is conducted: (1) direct inference using the weighted model, and (2) injecting the rule information into an LLM. For the latter, we systematically examine three levels of information: (i) providing only the rule, (ii) providing the rules with their weights, and (iii) providing the rules, weights and the linear model’s prediction as a reference. The layered design allows for a systematic evaluation of quality of rules and analysis of the trade-offs in accuracy and robustness for different ways of using rules. Using our framework, we can get high quality rules in natural languages, with optimal ways of utilizing them to perform rule based reasoning. The main contribution of this paper are as follows:
  1. 1.
    We propose RLIE, a unified paradigm for LLM-based probabilistic rule learning. It combines the natural language rule generation of LLMs with the global weighting and selection of regularized logistic regression to produce a self-consistent, collaborative and auditable rule set.
  2. 2.
    We design a hierarchical set of inference strategies to systematically compare direct inference with the linear combiner against three methods of injecting information into an LLM: rules-only, rules+weights and rules+weights+linear prediction reference. We provide empirical analysis and practical guidance regrading their respective trade-offs in terms of accuracy, robustness and calibration.
  3. 3.
    On multiple real-world datasets, RLIE achieves superior over all performance compared to a range of LLM-based methods. The learnt rule sets are more compact and semantically clearer, prompting knowledge discovery and human-AI consensus.
@@ 2.1 Classical Rule Learning and Its Applications
Rule Learning. Canonical approaches to rule learning include Inductive Logic Programming (ILP)(Cropper et al., 2022), which is centered on search and constraint optimization; association and pattern mining methods that rely on heuristics and pruning; and differentiable neuro-symbolic methods(Qiao et al., 2021; Glanois et al., 2022; Yang et al., 2024a) that couple symbolic rules with continuous representations. A common objective across these paradigms is to produce a set of human-readable and reusable discriminative rules from a structured candidate space. However, they are fundamentally dependent on a predefined predicate space.
Rule Organization and Reasoning. In the traditional paradigm, the interaction among rules is determined by their organization and the semantics of the inference process. One approach is the use of ordered rule lists or decision lists(Yang et al., 2017; Xu et al., 2024), which employ sequential matching to explicitly encode relationships of priority, exception, and default cases. This is well-suited for scenarios with clear sequential or override logic. Another approach involves unordered rule sets(Qiao et al., 2021; Yang & van Leeuwen, 2022), which support the parallel integration of multiple pieces of evidence, facilitating global trade-offs and model compression. Within the unordered setting, deterministic aggregation methods (e.g., logical OR/AND, thresholding, majority voting) are simple and intuitive but offer limited flexibility when faced with local conflicts, coverage gaps, or inter-rule dependencies. Consequently, many works have introduced probabilistic aggregation to model uncertainty and improve calibration. A common and interpretable implementation is the linear log-odds model, which treats the binary satisfaction of each rule $r_{j}(x)\in\{0,1\}$ on a sample $x$ as a feature and learns a weighted linear combination(Ruczinski et al., 2003):
$\mathrm{Pr}(y=1\mid x)=\sigma(\beta_{0}+\sum_{j}\beta_{j}r_{j}(x)).$
Our work builds upon this foundation. Traditional methods typically assume rules are expressed in a fixed, executable symbolic form, allowing for direct evaluation of their satisfaction. In contrast, our approach focuses on rules expressed in natural language, where satisfaction must be determined by an LLM. We adopt a two-level design: at the “local judgment” level, we leverage the expressive power and adaptability of LLMs to interpret individual rules; at the “global aggregation” level, we retain the robust and calibratable probabilistic linear combiner from the classical paradigm. This hybrid design captures the flexible semantics of natural language rules while ensuring the overall inference process remains transparent and robust.
@@ 2.2 LLM-based Rule Learning and Its Applications
Rule Learning. Large Language Models (LLMs) can directly propose hypotheses in natural language. A representative approach involves a “single-hypothesis refinement” loop, where the model maintains only the current best hypothesis and iteratively improves its quality and interpretability through feedback from validation sets or external executors(Qiu et al., 2023). In contrast, other methods adopt a “multi-hypothesis set” approach. For example, HypoGeniC(Zhou et al., 2024) generates multiple candidate hypotheses from data samples, maintains a top-$k$ set, and uses training signals to filter and update this set, ultimately forming an interpretable hypothesis bank. These two strategies exemplify the trade-off between deep optimization of a single hypothesis and the parallel retention of multiple diverse hypotheses.
Rule Organization and Reasoning. Rules expressed in natural language are typically reasoned over by the LLM that generated them. Existing work often passes one or more rules as context to an LLM, demonstrating that such information can enhance its reasoning capabilities(Zhang et al., 2024; Zhou et al., 2024). However, as the rules become more numerous or complex, the stability and instruction-following fidelity of LLMs can become unreliable. The black-box nature of LLMs makes it challenging to analyze the organization and interaction of these rules.
Our work is distinct from existing LLM-based rule learning approaches in that we are the first to explicitly combine LLMs with probabilistic methods to learn a set of weighted rules. This integration enables more principled sample selection during the iterative refinement stage and leads to more flexible and robust inference. To better analyze the synergy between this new rule paradigm and LLMs, we systematically compare direct inference using the probabilistic combiner with a hierarchy of strategies for injecting the rules, their weights, and the combiner’s predictions into an LLM. This allows us to provide empirical analysis and practical guidance on the trade-offs in accuracy and robustness.
@@ 3 Method
We consider a binary classification task on a labeled dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}$ is a natural language text and $y_{i}\in\{0,1\}$ is its corresponding label. The dataset $\mathcal{S}$ is partitioned into a training set $\mathcal{S}_{\mathrm{tr}}$, a validation set $\mathcal{S}_{\mathrm{val}}$ and a test set $\mathcal{S}_{\mathrm{te}}$. To perform rule based reasoning, our objective is to learn a set of natural language rules $\mathcal{H}^{\star}=\{h_{1},\ldots,h_{m}\}$ and a set of weights $\theta$ that aggregate the rules to model the discriminative relationship between $x$ and $y$. Here, $m\leq H$ where $H$ is a predefined capacity limit, and $h_{i}$ is a rule in natural language. Our method contains four stages, and the whole procedure is illustrated in Figure 1:
Figure 1: The pipeline of the RLIE framework. (1) Rule Generation: An LLM proposes an initial set of rules from a small sample of training data. (2) Logistic Regression: A regularized logistic regression model is trained on the full training set to learn probabilistic weights for the current rules. (3) Iterative Refinement: Prediction errors from the regression model are used to mine hard examples, which are then used to prompt the LLM to refine or generate new rules. This cycle continues until performance on a validation set converges. (4) Evaluation: The learnt rules as well as weights are evaluated with different inference strategies. Detailed explanations of different strategies are in Section 3.4.
  1. 1.
    Rule Generation: We initialize an empty rule set $\mathcal{H}:=\emptyset$. An LLM is prompted with a randomly selected subset of $\mathcal{S}_{\mathrm{tr}}$ to generate candidate rules, which are then filtered based on coverage to form the initial rule set $\mathcal{H}^{(1)}$.
  2. 2.
@@
    Logistic Regression: To obtain the optimal weights to combinatorially use the generated rules, logistic regression is performed to learn the rule weights $\theta=(\bm{\beta},b)$ on $\mathcal{S}_{\mathrm{tr}}$ with hyperparameters tuned using $\mathcal{S}_{\mathrm{val}}$.
  3. 3.
    Iterative Refinement: Hard examples are selected from $\mathcal{S}_{\mathrm{tr}}$ based on prediction errors. These samples, along with the current rule set $\mathcal{H}^{(t)}$, are fed back to the LLM to generate a set of new rules to revise existing rules. The union of newly generated rules and old rules can be directly used as $\mathcal{H}^{(t+1)}$, if it doesn’t exceed the capacity limit $H$, otherwise we need to discard some rules based on their classification accuracy. The rule weights $\theta^{(t+1)}$ are also updated. This process is repeatedly executed for several times to iteratively refine the rules, while the whole process is monitored on $\mathcal{S}_{\mathrm{val}}$ for early stopping.
  4. 4.
    Evaluation: After getting generated rules, we conduct a comprehensive evaluation on $\mathcal{S}_{\mathrm{te}}$ using two classes of inference strategies: direct prediction using the linear combiner, and injecting the learnt rules and their weights into an LLM for augmented reasoning, to evaluate the quality of rules with different inference strategies.
@@ 3.1 Rule Generation
Initialization and Candidate Generation. In the first iteration, we initialize the rule set as empty, $\mathcal{H}^{(0)}:=\emptyset$. We randomly sample $k$ examples from $\mathcal{S}_{\mathrm{tr}}$ and prompt an LLM to generate $h$ candidate rules (where $h\leq H$), denoted as $\mathcal{H}_{\mathrm{cand}}^{(1)}=\{{h}^{(1)}_{1},\ldots,{h}_{h}^{(1)}\}$.
Individual Rule Application. For any samples $x_{i}$ and candidate rule ${h}_{j}^{(1)}$, we prompt an LLM to obtain a local, ternary judgment
$z_{i,j}^{(1)}=\mathrm{LLM}\left(x_{i},{h}_{j}^{(i)}\right)\in\{-1,0,+1\},$
as $z_{i,j}$ means the evaluation results of applying the $j$-th rule on the $i$-th sample where +1+1 indicates a positive prediction, 1-1 a negative prediction and 0 signifies the rule is not applicable (“abstain”). Allowing abstention is crucial for explicitly modeling rule coverage, reducing misclassifications, and enabling sparse, robust combinations in the subsequent stage.
Coverage-based Filtering and Rule Set Update. For each candidate rule ${h}_{j}^{(1)}$, we compute its coverage on the training set as
$\mathrm{Cov}_{\mathrm{tr}}\left({h}_{j}^{(1)}\right)=\frac{1}{N_{\mathrm{tr}}}\sum_{i=1}^{N_{\mathrm{tr}}}\mathbb{I}\left[z_{i,j}^{(1)}\neq 0\right],$
and discard rules with coverage lower than a predefined threshold $\gamma$. After $\mathrm{Cov}_{\mathrm{tr}}$ based filtering, the valid rules generated here is denoted as $\mathcal{H}_{\mathrm{new}}^{(1)}$, and the rule set is updated as $\mathcal{H}^{(1)}=\mathcal{H}^{(0)}\cup\mathcal{H}_{\mathrm{new}}^{(1)}$.

### v2

@@ Abstract
Large Language Models (LLMs) can propose natural-language rules, circumventing the reliance on a predefined predicate space in traditional rule learning. However, existing LLM-based methods often neglect the global interactions among rules, and the potential of using fine-grained rule importance scores to calibrate neuro-symbolic reasoning remains underexplored. To address this gap, we introduce RLIE, a framework that integrates LLMs with probabilistic modeling to learn weighted rule sets in four stages: (1) Rule generation: proposing and filtering candidate rules via LLMs; (2) Logistic regression: learning sparse, calibrated weights for global rule selection; (3) Iterative refinement: revising the rule set with error-driven hard examples; and (4) Evaluation: validating the learned system via comparative inference paradigms. Across multiple real-world datasets and LLM backbones, our learned weighted rules achieve superior stability and accuracy, whereas rule-injection prompting yields mixed results and often degrades performance. These results suggest LLMs excel at semantic rule discovery but are less reliable at controlled probabilistic aggregation. Our findings highlight both the promise and the limits of LLMs for inductive reasoning, motivating a principled integration with classic probabilistic rule combination for reliable neuro-symbolic reasoning.
Machine Learning, ICML
@@ 1 Introduction
In many data-driven applications and scientific inquiry, the goal is increasingly shifting from pure prediction to constructing verifiable, reusable, and composable theories (Zhou et al., 2024; Yang et al., 2024a; Minh et al., 2022). Such theories enable explainable and auditable decisions and, in scientific contexts, can serve as building blocks for discovering underlying structures and accumulating knowledge (Yang et al., 2023, 2024b). These theories may be expressed as formal, structural statements (Cohen and others, 1995; Cropper and Morel, 2021) or as natural-language hypotheses [^1] In this paper, we use “rule” and “hypothesis” interchangeably, and refer to both as “rule”. (Zhou et al., 2024); regardless of form, they are declarative and testable units whose predictions can be verified by external evidence. A longstanding challenge is how to learn a collection of such testable hypotheses and integrate them into a coherent rule set that captures regularities, supports reliable decisions, and enables cumulative scientific understanding (Bazgir et al., 2025).
To ground this discussion, consider spam detection as a binary classification task. A predicate-based rule set might include three predicate statements:
1. $\mathrm{HasToken}(\texttt{prize},x)\wedge\mathrm{HasFile}(x)\rightarrow\mathrm{Spam}(x)$;
2. $\mathrm{InBlacklist}(x)\wedge\mathrm{HastLink}(x)\rightarrow\mathrm{Spam}(x)$;
3. $\mathrm{TitleAllCaps}(x)\wedge\mathrm{HasToken}(\texttt{free},x)\rightarrow\mathrm{Spam}(x)$.
A deterministic inference framework is to predict spam if any of the conditions is satisfied (Fürnkranz and Kliegr, 2015). Such classical rule learning methods depend on a predefined predicate space with limited expressiveness for unstructured inputs and open-ended semantics (Cerna and Cropper, 2024; Hocquette et al., 2024). In contrast, probabilistic approaches treat each rule’s firing as an indicator ($r_{j}\in\{0,1\}$) and learn global weights – e.g., via logistic regression – to combine them into a unified prediction (Ruczinski et al., 2003; Friedman and Popescu, 2008). Crucially, this probabilistic formulation allows for greater flexability, accommondating rules defined as either symbolic predicates or neural indicators.
Figure 1: The benefit of explicit aggregation. While direct LLM prompting struggles to resolve conflicts among rules with varying semantic levels (Top), RLIE’s structured approach—combining rule-wise judging with a weighted logistic combiner—effectively integrates evidence to yield a reliable prediction (Bottom).
The advent of large language models (LLMs) opens a new opportunity for learning expressive rules in natural languages: given a few examples and task context, LLMs can generate rules directly in natural language (Ellis, 2023; Singh et al., 2022). Such rules can articulate fine-grained conditions, relations, and exceptions that are cumbersome to enumerate with a fixed predicate space. However, existing LLM-based methods typically either iteratively optimize a single rule (Qiu et al., 2023) or generate multiple rules without considering their global interdependence. This often leads to misalignment between training and inference (Yang et al., 2023; Zhou et al., 2024; Yang et al., 2024b), leaving the challenge of learning a compact and coordinated rule set unresolved. More broadly, the probabilistic composition of natural-language rules remains under-studied: it is unclear whether LLMs can robustly apply rule weights for reliable aggregation. Motivated by our empirical findings on the limitations of LLM probabilistic reasoning (see Section 5.2), we propose a framework that combines the expressive power of LLMs with the calibrated composition of classical probabilistic models. See Figure 1 for an illustrative example.
To realize this synergy, we propose RLIE, a unified framework integrating Rule generation, Logistic regression, Iterative refinement, and Evaluation. RLIE first prompts an LLM to generate a pool of candidate rules. To enable learning, it uses the LLM to produce instance-level judgments for each rule. A regularized logistic regression model is then fitted to learn probabilistic weights, acting as explicit importance scores for rule selection. RLIE iteratively improves the rule set by mining hard examples to guide the LLM in revising or extending rules. Finally, the Evaluation stage verifies the system by comparing the learned logistic combiner against prompt-based inference. Our contributions can be summarized as follow:
1. We advocate a principled division of labor: LLMs handle semantic rule induction and local assessment, while statistical models provide globally calibrated composition, balancing semantic flexibility with inference robustness.
2. We propose a hard-example-driven refinement mechanism that improves the rule set through reflective revision and targeted extension, promoting rule quality and diversity.
3. We evaluate inference paradigms, finding that although LLMs excel at proposing interpretable rules, they struggle to effectively utilize rule weights for precise reasoning. This highlights a key deficiency in LLM reasoning and underscores the necessity of an external probabilistic combiner.
@@ 2.1 Classical Rule Learning and Its Applications
Rule Learning. Classical approaches encompass Inductive Logic Programming (ILP) (Cropper et al., 2022), association rule mining, and differentiable neuro-symbolic methods (Qiao et al., 2021; Glanois et al., 2022; Yang et al., 2024a). Despite their methodological differences, these paradigms share a common objective: inducing human-readable, reusable discriminative rules within a structured hypothesis space. A fundamental limitation remains their reliance on a predefined predicate vocabulary, which restricts expressiveness and hinders applicability to unstructured inputs with open-ended semantics.
Rule Organization and Reasoning. Inference semantics dictate how multiple rules interact to form a decision. One stream of research utilizes ordered rule lists (e.g., decision lists) to encode priority, exceptions, and defaults via sequential matching (Yang et al., 2017; Xu et al., 2024). Conversely, unordered rule sets integrate evidence in parallel, enabling global trade-offs and model compression (Qiao et al., 2021; Yang and van Leeuwen, 2022). Within unordered settings, deterministic aggregators (e.g., OR/AND, thresholding, voting) offer simplicity but often prove brittle under conflicts, coverage gaps, and inter-rule dependencies. Probabilistic aggregation mitigates these issues by modeling uncertainty and improving calibration. A prominent interpretable instantiation is the linear log-odds model, which treats rule satisfaction $r_{j}(x)\in\{0,1\}$ as a feature and learns a weighted combination (Ruczinski et al., 2003) $\Pr(y=1\mid x)=\sigma\!\left(\beta_{0}+\sum_{j}\beta_{j}r_{j}(x)\right).$
Our work builds upon this foundation but departs from the standard assumption that rules exist in a fixed, executable symbolic form with directly computable satisfactions. Instead, we focus on natural-language rules, whose instance-level satisfactions must be judged (and may involve abstention) by LLMs. This distinction motivates a principled division of labor: employing LLMs for local rule interpretation and instance-level judgments, while retaining a logistic combiner for global selection and calibrated aggregation. The resulting hybrid inference preserves the semantic expressiveness of natural-language rules while ensuring the stability and interpretability of the overall decision process.
@@ 2.2 LLM-based Rule Learning and Its Applications
Rule Learning. Current LLM-based induction falls into two paradigms: single-hypothesis refinement (Qiu et al., 2023), which iteratively optimizes one rule via feedback, and multi-hypothesis methods like HypoGeniC (Zhou et al., 2024) that maintain and update a candidate pool. However, these methods typically treat hypotheses in isolation, reflecting a trade-off between the intensive refinement of a single rule and the retention of a diverse rule set.
Rule Organization and Reasoning. Existing work predominantly uses the LLM as the inference engine via rule-augmented prompting (Zhang et al., 2024; Zhou et al., 2024). While effective in specific settings, implicit prompt-based aggregation becomes unstable as rule sets expand or conflict. Moreover, this black-box approach hinders error attribution, calibration, and the analysis of rule interactions, motivating our move toward explicit probabilistic composition.
In contrast, our work investigates explicit probabilistic composition for natural-language rules. We systematically compare direct probabilistic combination against prompt-based inference under progressively richer injected information (rules, weights, and prediction references), demonstrating the necessity of an external, calibrated aggregator.
@@ 3 Method
We consider a binary classification task on a labeled dataset $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}$ is a natural-language text and $y_{i}\in\{0,1\}$ is its label. We split $\mathcal{S}$ into training, validation, and test sets, denoted by $\mathcal{S}_{\mathrm{train}}$, $\mathcal{S}_{\mathrm{val}}$, and $\mathcal{S}_{\mathrm{test}}$, respectively. Our goal is to learn a set of natural-language rules $\mathcal{H}^{\star}=\{h_{1},\ldots,h_{m}\}$ together with aggregation weights $\theta$ that linearly combine these rules to model the discriminative relation between $x$ and $y$. We constrain the rule set size by $m\leq H$, where $H$ is a predefined capacity limit. Our framework consists of four stages (Figure 2):
Figure 2: Overview of the RLIE pipeline. (1) Rule Generation: An LLM proposes candidate rules from a small sample of $\mathcal{S}_{\mathrm{train}}$, which are then filtered by coverage to form an initial rule set. (2) Weight Learning: A regularized logistic regression combiner is trained on $\mathcal{S}_{\mathrm{train}}$ to learn rule weights. (3) Iterative Refinement: Hard samples are mined based on the combiner’s errors to prompt the LLM to revise or add rules; the rule set is pruned to satisfy a capacity limit and training proceeds with early stopping on $\mathcal{S}_{\mathrm{val}}$. (4) Evaluation: The learned rules and weights are assessed under different inference protocols.
1. Rule Generation: Starting from $\mathcal{H}:=\emptyset$, we prompt an LLM with a randomly sampled subset of $\mathcal{S}_{\mathrm{train}}$ to propose candidate rules, and filter them by coverage to obtain the initial rule set $\mathcal{H}^{(1)}$.
@@
2. Logistic Regression: Given $\mathcal{H}^{(t)}$, we train a logistic regression combiner on the whole training set $\mathcal{S}_{\mathrm{train}}$ to learn rule weights $\theta^{(t)}$, with hyperparameters selected on $\mathcal{S}_{\mathrm{val}}$.
3. Iterative Refinement: We select hard examples from $\mathcal{S}_{\mathrm{train}}$ with the highest prediction errors from the combiner, and feed them back to the LLM together with the current rule set $\mathcal{H}^{(t)}$ to generate new rules $\Delta\mathcal{H}^{(t)}$ to augment the rule set. We then merge the updated rules and enforce the capacity constraint: if $|\mathcal{H}^{(t)}\cup\Delta\mathcal{H}^{(t)}|>H$, we prune rules based on validation performance. The combiner is retrained to update $\theta^{(t+1)}$. We repeat this process with early stopping monitored on $\mathcal{S}_{\mathrm{val}}$.
4. Evaluation: We strictly evaluate the finalized rule set and weights on $\mathcal{S}_{\mathrm{test}}$. This stage quantifies the system’s reliability and compares the calibrated probabilistic inference (combiner) against uncalibrated LLM reasoning (prompting), verifying the effectiveness of weighted rule set.
Our prompts are provided in Appendix E.
@@ 3.1 Rule Generation
Initialization and Candidate Generation. We start with an empty rule set $\mathcal{H}^{(0)}:=\emptyset$. In the first iteration, we uniformly sample $k$ examples from $\mathcal{S}_{\mathrm{train}}$ and prompt an LLM to propose $h$ candidate natural-language rules ($h\leq H$), denoted by $\mathcal{H}_{\mathrm{cand}}^{(1)}=\{h^{(1)}_{1},\ldots,h^{(1)}_{h}\}$.
Individual Rule Application. For each candidate rule $h_{j}^{(1)}$ and sample $x_{i}$ from the whole training set, we query an LLM-based judge to produce a ternary judgment $z_{i,j}^{(1)}=\mathrm{LLM}\!\left(x_{i},\,h_{j}^{(1)}\right)\in\{-1,0,+1\},$ where +1+1 and 1-1 indicate predictions for the positive and negative class, respectively, and 0 denotes abstention when the LLM deems the rule inapplicable to sample $x_{i}$. This explicitly models rule coverage and mitigates spurious forced predictions, facilitating sparse and robust aggregation in the subsequent stage.
Coverage-based Filtering and Rule Set Update. We measure the coverage of each candidate rule $h_{j}^{(1)}$ on the training set as: $\mathrm{Cov}_{\mathrm{train}}\!\left(h_{j}^{(1)}\right)=\frac{1}{N_{\mathrm{train}}}\sum_{i=1}^{N_{\mathrm{train}}}\mathbb{I}\!\left[z_{i,j}^{(1)}\neq 0\right].$ We discard rules from $\mathcal{H}_{\mathrm{cand}}^{(1)}$ with $\mathrm{Cov}_{\mathrm{train}}\!\left(h_{j}^{(1)}\right)<\gamma$ to ensure that the rule set is general and applicable to a significant portion of samples. We denote the remaining rules by $\mathcal{H}_{\mathrm{new}}^{(1)}$. The rule set is then updated as $\mathcal{H}^{(1)}=\mathcal{H}^{(0)}\cup\mathcal{H}_{\mathrm{new}}^{(1)}$.